### v1

@@ Abstract
Empirical studies using Regression Discontinuity (RD) designs often explore heterogeneous treatment effects based on pretreatment covariates. However, the lack of formal statistical methods has led to the widespread use of ad hoc approaches in applications. Motivated by common empirical practice, we develop a unified, theoretically grounded framework for RD heterogeneity analysis. We show that a fully interacted local linear (in functional parameters) model effectively captures heterogeneity while still being tractable and interpretable in applications. The model structure holds without loss of generality for discrete covariates, while for continuous covariates our proposed (local functional linear-in-parameters) model can be potentially restrictive, but it nonetheless naturally matches standard empirical practice and offers a causal interpretation for RD applications. We establish principled bandwidth selection and robust bias-corrected inference methods to analyze heterogeneous treatment effects and test group differences. We provide companion software to facilitate implementation of our results. An empirical application illustrates the practical relevance of our methods.
@@ 1 Introduction
Analyzing heterogeneous treatment effects based on pretreatment covariates is crucial in modern causal inference. While the average causal effect provides a broad measure of a treatment’s effectiveness and informs general policy decisions, uncovering covariate heterogeneity patterns is essential for evaluating substantive hypotheses and specific policy interventions. Covariate-based heterogeneous treatment effects can help understand fairness and differential impacts across subpopulations, develop targeted interventions, and optimize social policies.
@@
Over the past decade, the regression discontinuity (RD) design has become a prominent tool for causal inference, leading to extensive methodological advancements for identification, estimation, and inference for average treatment effects in various contexts. However, despite its widespread adoption, there is a notable lack of rigorous methods for studying heterogeneity based on pretreatment covariates, resulting in many ad-hoc empirical approaches. We address this gap by developing practical, theoretically-grounded tools for heterogeneity analysis in RD designs, putting forward a unified methodology for empirical researchers studying heterogeneous treatment effects based on pretreatment covariates.
@@
Extending RD estimation and inference to encompass heterogeneous treatment effects presents a high-dimensional, nonparametric challenge: while theoretically feasible, it is often impractical without model restrictions. Consequently, empirical researchers typically use semiparametric models, and thus explore heterogeneity in limited dimensions. Following empirical practice, we study the most common approach for RD treatment effect heterogeneity analysis: local least squares regression with linear-in-parameters interactions with pretreatment covariates. This model balances practical applicability with the flexible (nonparametric) nature of the RD design.
@@
Our first contribution is to clarify the conditions under which a local regression with linear interactions accurately recovers meaningful causal heterogeneous RD treatment effects. We show that heterogeneous effects are identifiable when potential outcomes follow a local linear-in-parameters functional coefficient model in the pretreatment variables. This structure is both flexible and interpretable, and is without loss of generality for binary (and orthogonal) covariates. Without this structure, the linear interactions model leads to a less interpretable and useful probability limit, as the nonparametric nature of the estimation method hampers its usual causal interpretation emerging from the best linear prediction viewpoint. The supplemental appendix computes the parameters emerging as the probability limit of a local linear regression with linear interactions under general misspecification, and demonstrates by example that a large smoothing bias emerges without our proposed identifying assumption. Therefore, from a practical perspective, we show that (i) heterogeneity analysis based on local polynomial regression methods with interactions is fully warranted without additional assumptions when the pretreatment covariates are binary orthogonal variables (e.g., indicator variables for each mutually exclusive level of a categorical variable), but (ii) heterogeneity analysis requires a semiparametric identifying assumption in the general case of discrete or continuous covariates for causal inference.
@@
We then present formal methods for estimation and inference of heterogeneous RD treatment effects, developing optimal bandwidth selection, point estimation, and robust bias-corrected inference. Our key large-sample econometric results include consistency, a mean squared error expansion, a central limit theorem, and robust standard error estimators to both heteroskedasticity and clustering. We show that the optimal bandwidth depends on the specific inference target, implying different optimal bandwidths for average effects, subgroups, and group difference tests. Our goal is to standardize empirical practice, where bandwidth selection is often ad-hoc, particularly for heterogeneity analysis. From that perspective, we identify conditions under which a single bandwidth can be validly used for the entire analysis, thereby enhancing the practical appeal of our proposed methods.
@@
Despite its major role in empirical work, covariate adjustment for heterogeneity analysis in RD designs has been surprisingly less explored in the literature than needed. The only antecedents we are aware of are Hsu and Shen (2019), who use moment condition methods to test for the presence of covariate-based heterogeneity, and Reguly (2021) and Alcantara et al. (2025), who leverage machine learning methods to recover nonparametric covariate-based heterogeneity. None of these contributions study the properties of the most common approach in empirical work: semi-linear local polynomial regression with pretreatment covariate interactions. Our paper addresses this gap in the literature by setting up a standard, most tractable local least squares framework with covariate-interactions, and then providing novel identification, estimation, and robust bias-corrected inference methods to support past, current, and future heterogeneity analysis in RD applications.
@@
By introducing a comprehensive framework for identification, estimation, and inference in RD designs, we aim to unify the study of heterogeneous treatment effects in empirical research. The absence of such a framework has led to a variety of empirical approaches—some valid, others not—creating inconsistencies and confusion in the literature. To illustrate this issue, Section SA3 in the supplemental appendix compiles a sample of papers from AEA journals over the past decade that analyze covariate heterogeneity in RD settings, provided they offer sufficient methodological details. While not exhaustive or necessarily representative, this review reveals three common patterns. First, most studies examine heterogeneity for binary covariates, but often by discretizing categorical or continuous variables. Second, a single bandwidth, and often selected in an ad hoc manner, is typically used for both average and heterogeneous treatment effects analysis. Third, and most importantly, inference is rarely conducted properly, regardless of bandwidth choice or covariate granularity. Our proposed methods maintain the simplicity and interpretability of standard empirical practice while introducing objective procedures that bring validity and rigor.
@@
We illustrate our methodological results by revisiting the analysis in Akhtari et al. (2022), who study how political turnover in mayoral elections in Brazil affects public service provision by local governments. The analysis exploits close elections, a canonical RD setting, using a rich data set of Brazilian mayoral elections. Klašnja and Titiunik (2017) originally analyzed this data and RD setting, documenting a negative incumbency advantage; see further references therein. We begin by replicating the results in Akhtari et al. (2022), who considered both standard RD treatment effect and a limited measure of covariate-heterogeneity based on municipality-level monthly household income. We recover their point estimates of heterogeneity for income dichotomized as above or below median, and also report valid uncertainty quantification; our results are qualitatively in line with theirs. Then, we further leverage our framework and methods to study income discretized by quartile and by decile, revealing further nuance in the heterogeneity of treatment effects by income level. To conclude, we also show that using income directly (as a continuous variable) provides an accurate summary of the heterogeneity in this application, which is (i) highly compatible with the discretized estimates, (ii) directly interpretable, and (iii) more efficient for estimation and inference. Specifically, we consider two local linear-in-parameters semiparametric heterogeneity models: one taking income linearly, while the other taking income quadratically. This empirical exercise demonstrates how our methods can be used to enhance RD heterogeneity analysis, while also highlighting the importance of having a clear understanding of how best to capture heterogeneity in the econometric model.
@@
Our paper contributes to the methodological literature on RD designs (see Cattaneo and Titiunik, 2022, for a review). In particular, it sheds new light on the role of pretreatment covariates for identification, estimation, and inference of heterogeneous RD treatment effects. Cattaneo et al. (2023) overviews covariate adjustment in RD designs and discusses its three most useful roles: (i) increasing estimation efficiency, (ii) changing the parameter of interest, and (iii) heterogeneity analysis. For efficiency gains, Calonico et al. (2019), Arai et al. (2025), Ma and Yu (2024), and references therein, give results leveraging pretreatment covariates to improve precision in the estimation of the average RD treatment effect. There is also a rich literature on the inclusion of covariates for identification of other RD treatment effects: for example, see Angrist and Rokkanen (2015) and Cattaneo et al. (2021) for extrapolation, Peng and Ning (2021) and Caetano et al. (2024) for weighted average treatment effects, Grembi et al. (2016) for difference-in-discontinuities, and Frölich and Huber (2019) for nonparametric methods, among many others. As mentioned above, the role of covariates for heterogeneity analysis has received the least attention in the literature, and our paper is the first to study it in the context of local semi-linear regression with interactions, the most common approach used in practice. Finally, it is worth mentioning that covariates are sometimes used in an attempt to “fix” a “broken” RD design. This is not valid without strong assumptions: it is analogous to employ covariate adjustment to “fix” a “broken” randomized assignment in experimental settings. Our paper does not speak to this practice, and we recommend against it. See Cattaneo et al. (2023) for further discussion.
@@
The rest of the paper proceeds as follows. Section 2 gives details on the RD setup and estimation. Our main technical results are summarized in Section 3. We then present an empirical illustration of our proposed methods in Section 4. Section 5 concludes. The supplemental appendix contains more general, technical results, all proofs, implementation details, and additional numerical results. The companion software package rdhte offers general-purpose implementation of our methodological results, and is available at https://rdpackages.github.io/rdhte/. See Calonico et al. (2025) for details.
@@ 2 Setup
We consider the standard sharp regression discontinuity design setting with a continuous running variable (see Cattaneo et al., 2020, 2024, for a practical introduction). The defining characteristic is that a binary treatment $T_{i}\in\{0,1\}$ is assigned to each unit based on whether the continuous running variable $X_{i}$ is above or below a known cutoff, denoted by $c$. Thus, $T_{i}=\mathbbm{1}\{X_{i}\geq c\}$. The observed outcome is $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$, where $Y_{i}(t)$ is the potential outcome for unit $i$ under treatment $T_{i}=t$. In addition to $Y_{i}$, $T_{i}$, and $X_{i}$, we observe a vector $\mathbf{W}_{i}\in\operatorname{\mathbb{R}}^{d}$ of pretreatment covariates to be used for heterogeneity analysis. Distinct from $\mathbf{W}_{i}$, we may have an additional set of covariates $\mathbf{Z}_{i}$ used for efficiency purposes only; see below. We assume a random sample $(Y_{i},T_{i},X_{i},\mathbf{W}_{i},\mathbf{Z}_{i})$, for $i=1,\dots,n$, is observed.
The usual parameter of interest in sharp RD is the average effect of $T_{i}$ on $Y_{i}$ at the cutoff. To simplify notation, we set the cutoff to be $c=0$. Then, the average RD treatment effect is
@@
Identification of $\tau$ follows from standard continuity assumptions (Hahn et al., 2001), while estimation and inference is typically done with local linear regression and robust bias correction (Calonico et al., 2014). More precisely, the point estimate $\dot{\tau}$ of the causal parameter $\tau$ is the coefficient on $T_{i}$ in a least squares regression of $Y_{i}$ on a constant, $T_{i},X_{i}$, and $T_{i}\cdot X_{i}$, using only the observations such hat $|X_{i}-0|\leq h$, for a bandwidth $h$, and weighted by $K((X_{i}-0)/h)$, for a kernel function $K(\cdot)$. This fit, and those below, are defined formally in the Supplemental Appendix (Section SA1). We denote the resulting estimated equation by
@@
The coefficient $\dot{\tau}$ is identical to the difference in intercepts from fitting separate (local, weighted) least squares regressions on either side of the cutoff. Conceptually, it is important to remember that (2.2) is formally a nonparametric local polynomial regression with the polynomial degree set to one. Finally, although point estimation is done with (local, weighted) linear models, we stress that inference is then conducted using robust bias correction, which we adapt for heterogeneity below.
Calonico et al. (2019) studied the addition of the pretreatment covariates to this model, but aiming exclusively for more efficient estimation of the RD average treatment effect $\tau$, not for heterogeneity analysis. They showed that efficiency gains could be achieved by including covariates linearly, but not interacting them with the treatment or running variable. Specifically, for pretreatment covariates $\mathbf{Z}_{i}$ (used for efficiency only and hence necessarily distinct from $\mathbf{W}_{i}$), Calonico et al. (2019) studied $\widetilde{\tau}$ from the (local, weighted) least squares fit given by
@@
Unlike $\dot{\tau}$, the estimated coefficient $\widetilde{\tau}$ cannot be obtained from two separate fits because of the common coefficient on $\mathbf{Z}_{i}$. Despite being linear in both the running variable and covariates, it is important to note that (2.3) is nonparametric in $X_{i}$ (local polynomial of degree one, localized using the bandwidth $h$) but it is a linear projection in $\mathbf{Z}_{i}$, and thus the interpretation of the estimated coefficient $\widetilde{\bm{\gamma}}$ is only as a best linear projection (unless the population mean function happens to be linear). That is, the nonparametric smoothing is only on the running variable. This respects the local nature of the RD design while being simple and tractable in applications. Calonico et al. (2019) proved that, nonetheless, $\widetilde{\tau}$ recovers the average effect $\tau$, characterized conditions for efficiency gains, and provided optimal bandwidth selection and valid inference using bias robust bias correction. Ma and Yu (2024) strengthened the efficiency argument for this regression fit. The exclusion of interactions with the covariates in (2.3) is crucial for correct estimation of the average effect $\tau$ of (2.1), but by construction does not allow for learning heterogeneity. Because our present goal is explicitly to learn about heterogeneous treatment effects, we need to allow some form of interaction between the pretreatment covariates and treatment indicator.
@@
We consider the local weighted least squares regression with full interactions between the treatment indicator and the pretreatment covariates, which may be discrete or continuous. That is, our main task is analyzing the interpretation and properties of the regression fit:
@@
As with Equations (2.2) and (2.3), the fit (2.5) is a local polynomial (of degree 1) approximation in $X_{i}$, with bandwidth $h$ and kernel weights $K(X_{i}/h)$ around the cutoff $c=0$. In this case, this is true for the main effects as well as the interaction terms: while $\widetilde{\bm{\gamma}}$ in (2.3) is a linear projection coefficient with no intended approximation power (or causal interpretation), the semiparametric fit (2.5) delivers local polynomial approximations to unknown functions of $X_{i}$ only, on either side of the cutoff, that form linear interactions with $\mathbf{W}_{i}$. This is formalized in Assumption 2 below and in the supplemental appendix in full generality (Section SA1.3). Importantly, such an approximation cannot consistently estimate an arbitrary $\kappa(\mathbf{w})$; its causal interpretation crucially relies on Assumption 2 below. To reiterate, this is true because the regression (2.5) is nonparametric in $X_{i}$ but not $\mathbf{W}_{i}$; there is no approximation to unknown functions of $\mathbf{W}_{i}$.
@@
In applications, the vector $\mathbf{W}_{i}$ must therefore be specified by the researcher to capture the relevant heterogeneity. Changing the covariates, or their specification, included in the regression changes the target estimand and its interpretation. With discrete variables, the vector $\mathbf{W}_{i}$ should contain indicators for every relevant category, unless the researcher decides ex-ante that certain categories are not to be studied. For example, if two binary variables $(B_{i,1},B_{i,2})$ are present, then $\mathbf{W}_{i}=\big{(}(1-B_{i,1})B_{i,2},\ B_{i,1}(1-B_{i,2}),\ B_{i,1}B_{i,2}\big{)}^{\prime}$, with $B_{i,1}=B_{i,2}=0$ being the baseline category. A multi-level discrete variable would require that $\mathbf{W}_{i}$ contain an indicator for each level beyond the baseline, which forms a collection of orthogonal binary variables. For example, if $\mathcal{W}$ is the finite set of values the covariate of interest $W_{i}$ takes (e.g., years of education), then $\mathbf{W}_{i}=(\mathbbm{1}(W_{i}=w),w\in\mathcal{W})^{\prime}$. In all of these cases, the structure of Assumption 2 holds without loss of generality. Continuous variables can be included directly into $\mathbf{W}_{i}$, in which case $\widehat{\bm{\xi}}$ have the standard interpretation of a slope coefficient in a linear model. Further, transformations of continuous variables (such as polynomials) can be added for additional flexibility. Alternatively, the researcher might estimate the heterogeneous treatment effects one variable at a time, which may not capture the full heterogeneity pattern but may provide a simple and intuitive summary of different dimensions of heterogeneity. Section 4 illustrates different approaches and interpretations empirically. (See Calonico et al. (2025) for further examples and discussion.)
@@
As with (2.2), and in contrast with (2.3), the coefficients in (2.5) can be obtained by first fitting separate regressions on either side of the cutoff, and then taking appropriate differences. It is also possible to run separate models for different subgroups of discrete variables via (2.2) to recover the estimates in (2.5), but the joint estimation procedure is simple and intuitive and, thus, appealing for practice. Joint estimation and inference also simplify concerns of bandwidth selection, which otherwise may change depending on the group of interest or the hypothesis being tested. Furthermore, joint estimation is also natural if additional controls are added: following Calonico et al. (2019), it is possible to include some covariates for efficiency gains, while other covariates for heterogeneity analysis, thereby obtaining more efficient estimation and inference for $\kappa(\mathbf{w})$. This mixed covariate adjustment approach augments (2.5) with additional covariates $\mathbf{Z}_{i}$ that are not interacted with $T_{i}$ nor $X_{i}$, but interacted with $\mathbf{W}_{i}$. That is, one would fit

### v3

@@ Abstract
Empirical studies using Regression Discontinuity (RD) designs often explore heterogeneous treatment effects based on pretreatment covariates, even though no formal statistical methods exist for such analyses. This has led to the widespread use of ad hoc approaches in applications. Motivated by common empirical practice, we develop a unified, theoretically grounded framework for RD heterogeneity analysis. We show that a fully interacted local linear (in functional parameters) model effectively captures heterogeneity while still being tractable and interpretable in applications. The model structure holds without loss of generality for discrete covariates. Although our proposed model is potentially restrictive for continuous covariates, it naturally aligns with standard empirical practice and offers a causal interpretation for RD applications. We establish principled bandwidth selection and robust bias-corrected inference methods to analyze heterogeneous treatment effects and test group differences. We provide companion software to facilitate implementation of our results. An empirical application illustrates the practical relevance of our methods.
@@ 1 Introduction
Analyzing heterogeneous treatment effects based on pretreatment covariates is crucial in modern causal inference. While the average causal effect provides a broad measure of a treatment’s effectiveness and informs general policy decisions, uncovering patterns of covariate heterogeneity is often essential for evaluating substantive hypotheses and specific policy interventions. Covariate-based heterogeneous treatment effects can help understand fairness and differential impacts across subpopulations, develop targeted interventions, and optimize social policies.
@@
Over the past decade, the regression discontinuity (RD) design has become a prominent tool for causal inference, leading to extensive methodological advancements for identification, estimation, and inference for average treatment effects in various contexts. Despite its widespread adoption, there is still a notable lack of rigorous methods for studying heterogeneity based on pretreatment covariates, resulting in many ad-hoc empirical approaches. We address this gap by developing practical, theoretically-grounded tools for heterogeneity analysis in RD designs, putting forward a unified methodology for empirical researchers studying heterogeneous treatment effects based on pretreatment covariates.
@@
Extending RD estimation and inference to encompass heterogeneous treatment effects presents a high-dimensional, nonparametric challenge: while theoretically feasible, it is often impractical without model restrictions. Consequently, empirical researchers typically use semiparametric models and explore heterogeneity in limited dimensions. Following empirical practice, we study the most common approach for RD treatment effect heterogeneity analysis: local least squares regression with linear-in-parameters interactions with pretreatment covariates. This model balances practical applicability with the flexible, nonparametric nature of the RD design.
@@
Our first contribution is to clarify the conditions under which a local regression with linear interactions accurately recovers meaningful causal heterogeneous RD treatment effects. We show that heterogeneous effects are identifiable when potential outcomes follow a local linear-in-parameters functional coefficient model in the pretreatment variables. This structure is both flexible and interpretable, and is without loss of generality for binary (and orthogonal) covariates. Without this structure, the linear interactions model leads to a less interpretable and less useful probability limit, as the nonparametric nature of the estimation method hampers its usual causal interpretation emerging from the best linear prediction viewpoint. Therefore, from a practical perspective, we show that (i) heterogeneity analysis based on local polynomial regression methods with interactions is fully warranted without additional assumptions when the pretreatment covariates are binary orthogonal variables (e.g., indicator variables for each mutually exclusive level of a categorical variable), but (ii) heterogeneity analysis requires a semiparametric identifying assumption in the general case of discrete or continuous covariates for causal inference.
@@
We then present formal methods for estimation and inference of heterogeneous RD treatment effects, developing optimal bandwidth selection, point estimation, and robust bias-corrected inference. Our key large-sample econometric results include consistency, a mean squared error expansion, a central limit theorem, and standard error estimators robust to both heteroskedasticity and clustering. We show that the optimal bandwidth depends on the specific inference target, implying different optimal bandwidths for average effects, subgroups, and group difference tests. Our goal is to standardize empirical practice, where bandwidth selection is often ad-hoc, particularly for heterogeneity analysis. From that perspective, we identify conditions under which a single bandwidth can be validly used for the entire analysis, thereby enhancing the practical appeal of our proposed methods.
@@
Despite its major role in empirical work, covariate adjustment for heterogeneity analysis in RD designs has been minimally explored. The only antecedents we are aware of are Hsu and Shen (2019, 2021), who use moment condition methods to test for the presence of covariate-based heterogeneity, and Reguly (2021) and Alcantara et al. (2025), who leverage machine learning methods to recover nonparametric covariate-based heterogeneity. None of these contributions study the properties of the most common approach in empirical work: semi-linear local polynomial regression with pretreatment covariate interactions. We address this gap by setting up a standard, tractable local least squares framework with covariate interactions, and then providing novel identification, estimation, and robust bias-corrected inference methods to support heterogeneity analysis in RD applications.
@@
By introducing a comprehensive framework for identification, estimation, and inference in RD designs, we aim to unify the study of heterogeneous treatment effects in empirical research. The absence of such a framework has led to a variety of empirical approaches—some valid, others not—creating inconsistencies and confusion in the literature. To illustrate this issue, Section SA3 in the supplemental appendix compiles a sample of papers from AEA journals over the past decade that analyze covariate heterogeneity in RD settings, provided they offer sufficient methodological details. While not exhaustive or necessarily representative, this review reveals three common patterns. First, most studies examine heterogeneity for binary covariates, but often by discretizing categorical or continuous variables. Second, a single bandwidth, often selected in an ad hoc manner, is typically used for both average and heterogeneous treatment effects analysis. Third, and most importantly, inference is rarely conducted properly, regardless of bandwidth choice or covariate granularity. Our proposed methods maintain the simplicity and interpretability of standard empirical practice while introducing objective procedures that bring validity and rigor.
@@
We illustrate our methodological results by revisiting the analysis in Akhtari et al. (2022), who study how political turnover in mayoral elections in Brazil affects public service provision. The analysis exploits close elections, a canonical RD setting, using a rich data set of Brazilian mayoral elections. Klašnja and Titiunik (2017) originally analyzed these data and RD setting, documenting a negative incumbency advantage; see further references therein. We begin by replicating the results in Akhtari et al. (2022), who considered both standard RD treatment effects and a limited measure of covariate-heterogeneity based on municipality-level monthly household income. We recover their point estimates of heterogeneity for income dichotomized as above or below the median, and also report valid uncertainty quantification; our results are qualitatively in line with theirs. Then, we use our framework to study income discretized by quartile and by decile, revealing further nuance in the heterogeneity of treatment effects by income level. To conclude, we also show that using income directly (as a continuous variable) provides an accurate summary of the heterogeneity in this application, which is (i) highly compatible with the discretized estimates, (ii) directly interpretable, and (iii) more efficient for estimation and inference. For implementation, we consider two local linear-in-parameters semiparametric heterogeneity models: one taking income linearly, and the other quadratically. This empirical exercise demonstrates how our methods can be used to enhance RD heterogeneity analysis and highlights the importance of clearly understanding how to best capture heterogeneity in the econometric model.
@@
Our paper contributes to the methodological literature on RD designs (see Cattaneo and Titiunik, 2022, for a review), shedding new light on the role of pretreatment covariates for identification, estimation, and inference of heterogeneous RD treatment effects. Cattaneo et al. (2023) overviews covariate adjustment in RD designs and discusses its three most useful roles: (i) increasing estimation efficiency, (ii) changing the parameter of interest, and (iii) heterogeneity analysis. For efficiency gains, Calonico et al. (2019), Ma and Yu (2024), Arai et al. (2025), and references therein, give results leveraging pretreatment covariates to improve precision in the estimation of the average RD treatment effect. There is also a rich literature on the inclusion of covariates for identification of other RD treatment effects: for example, see Angrist and Rokkanen (2015) and Cattaneo et al. (2021) for extrapolation, Peng and Ning (2021) and Caetano et al. (2024) for weighted average treatment effects, Grembi et al. (2016) for difference-in-discontinuities, and Frölich and Huber (2019) for nonparametric methods, among many others. As mentioned, the role of covariates for heterogeneity analysis has received the least attention in the literature, and our paper is the first to study it in the context of local semi-linear regression with interactions, the most common approach used in practice. Finally, covariates are sometimes used in an attempt to “fix” a “broken” RD design. This is not valid without strong assumptions: it is analogous to employ covariate adjustment to “fix” a “broken” randomized assignment in experimental settings. Our paper does not speak to this practice, and we recommend against it. See Cattaneo et al. (2023) for further discussion.
@@
The rest of the paper proceeds as follows. Section 2 gives details on the RD setup and estimation. Our main technical results are summarized in Section 3, and we present an empirical illustration in Section 4. Section 5 concludes. The supplemental appendix contains more general, technical results, all proofs, implementation details, and additional numerical results. The software package rdhte (Calonico et al., 2025) provides a general-purpose implementation of our methodological results and is available at https://rdpackages.github.io/rdhte/.
@@ 2 Setup
We consider the standard sharp regression discontinuity design setting with a continuous running variable (see Cattaneo et al., 2020, 2024, for a practical introduction). The defining characteristic is that a binary treatment $T_{i}\in\{0,1\}$ is assigned to each unit based on whether the continuous running variable $X_{i}$ is above or below a known cutoff, denoted by $c$. Thus, $T_{i}=\mathbbm{1}\{X_{i}\geq c\}$. The observed outcome is $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$, where $Y_{i}(t)$ is the potential outcome for unit $i$ under treatment $T_{i}=t$. In addition to $Y_{i}$, $T_{i}$, and $X_{i}$, we observe a vector $\mathbf{W}_{i}\in\operatorname{\mathbb{R}}^{d}$ of pretreatment covariates to be used for heterogeneity analysis. Distinct from $\mathbf{W}_{i}$, we may have an additional set of covariates $\mathbf{Z}_{i}\in\operatorname{\mathbb{R}}^{d_{z}}$ used for efficiency purposes only; see below. We assume a random sample $(Y_{i},T_{i},X_{i},\mathbf{W}_{i},\mathbf{Z}_{i})$, for $i=1,\dots,n$, is observed.
The usual parameter of interest in sharp RD designs is the average effect of $T_{i}$ on $Y_{i}$ at the cutoff. To simplify notation, we set the cutoff to be $c=0$. Then, the average RD treatment effect is
@@
Identification of $\tau$ follows from standard continuity assumptions (Hahn et al., 2001), while estimation and inference are typically done with local linear regression and robust bias correction (Calonico et al., 2014). More precisely, the point estimate $\dot{\tau}$ of the causal parameter $\tau$ is the coefficient on $T_{i}$ in a least squares regression of $Y_{i}$ on a constant, $T_{i},X_{i}$, and $T_{i}\cdot X_{i}$, using only the observations such that $|X_{i}-0|\leq h$, for a bandwidth $h$, and weighted by $K((X_{i}-0)/h)$, for a kernel function $K(\cdot)$. This fit, and those below, are defined formally in the Supplemental Appendix (Section SA1.2). We denote the resulting estimated equation by
@@
The coefficient $\dot{\tau}$ is identical to the difference in intercepts from fitting separate (local weighted) least squares regressions on either side of the cutoff. Conceptually, it is important to remember that (2.2) is formally a nonparametric local polynomial regression with the polynomial degree set to one. Finally, although point estimation is done with (local weighted) linear models, we stress that inference is then conducted using robust bias correction, which we adapt for heterogeneity below.
Calonico et al. (2019) studied the addition of the pretreatment covariates to this model, but aiming exclusively for more efficient estimation of the RD average treatment effect $\tau$, not for heterogeneity analysis. They showed that efficiency gains could be achieved by including covariates linearly, but not interacting them with the treatment or running variable. Specifically, for pretreatment covariates $\mathbf{Z}_{i}$ (used for efficiency only and hence necessarily distinct from $\mathbf{W}_{i}$), Calonico et al. (2019) studied $\widetilde{\tau}$ from the (local weighted) least squares fit given by
@@
Unlike $\dot{\tau}$, the estimated coefficient $\widetilde{\tau}$ cannot be obtained from two separate fits because of the common coefficient on $\mathbf{Z}_{i}$. Despite being linear in both the running variable and covariates, equation (2.3) is nonparametric in $X_{i}$ (local polynomial of degree one, localized using the bandwidth $h$) but it is a linear projection in $\mathbf{Z}_{i}$, and thus the interpretation of the estimated coefficient $\widetilde{\bm{\gamma}}$ is only as a best linear projection (unless the population mean function happens to be linear). That is, the nonparametric smoothing is only on the running variable. This respects the local nature of the RD design while being simple and tractable in applications. Calonico et al. (2019) proved that $\widetilde{\tau}$ recovers the average effect $\tau$, characterized conditions for efficiency gains, and provided optimal bandwidth selection and valid inference using bias robust bias correction. Ma and Yu (2024) strengthened the efficiency argument for this regression fit. The exclusion of interactions with the covariates in (2.3) is crucial for correct estimation of the average effect $\tau$ of (2.1), but by construction does not allow for learning heterogeneity. Because our present goal is explicitly to learn about heterogeneous treatment effects, we need to allow some form of interaction between the pretreatment covariates and treatment indicator.
@@
We consider the local weighted least squares regression with full interactions between the treatment indicator and the pretreatment covariates, which may be discrete or continuous. Our main task is analyzing the interpretation and properties of the regression fit
@@
As with Equations (2.2) and (2.3), the fit (2.5) is a local polynomial (of degree 1) approximation in $X_{i}$, with bandwidth $h$ and kernel weights $K(X_{i}/h)$ around the cutoff $c=0$. In this case, this is true for the main effects as well as the interaction terms: while $\widetilde{\bm{\gamma}}$ in (2.3) is a linear projection coefficient with no intended approximation power (or causal interpretation), the semiparametric fit (2.5) delivers local polynomial approximations to unknown functions of $X_{i}$, on either side of the cutoff, that form linear interactions with $\mathbf{W}_{i}$. This is formalized in Assumption 2 below and in the supplemental appendix in full generality (see Section SA1.3). Importantly, such an approximation cannot consistently estimate an arbitrary $\kappa(\mathbf{w})$; its causal interpretation crucially relies on Assumption 2. This is true because the regression (2.5) is nonparametric in $X_{i}$ but not in $\mathbf{W}_{i}$; there is no approximation to unknown functions of $\mathbf{W}_{i}$.
@@
In applications, the vector $\mathbf{W}_{i}$ must therefore be specified by the researcher to capture the relevant heterogeneity. Changing the covariates included in the regression, or their specification, changes the target estimand and its interpretation. With discrete variables, the vector $\mathbf{W}_{i}$ should contain indicators for every relevant category, unless the researcher decides ex-ante that certain categories are not to be studied. For example, if two binary variables $(B_{i,1},B_{i,2})$ are present, then $\mathbf{W}_{i}=\big{(}(1-B_{i,1})B_{i,2},\ B_{i,1}(1-B_{i,2}),\ B_{i,1}B_{i,2}\big{)}^{\prime}$, with $B_{i,1}=B_{i,2}=0$ being the baseline category. A multi-level discrete variable would require that $\mathbf{W}_{i}$ contain an indicator for each level beyond the baseline, which forms a collection of orthogonal binary variables. For example, if $\mathcal{W}$ is the finite set of values taken by the covariate of interest $W_{i}$ (e.g., years of education), then $\mathbf{W}_{i}=(\mathbbm{1}(W_{i}=w),w\in\mathcal{W})^{\prime}$. In all of these cases, the structure of Assumption 2 holds without loss of generality. Continuous variables can be included directly into $\mathbf{W}_{i}$, in which case $\widehat{\bm{\xi}}$ has the standard interpretation of a slope coefficient in a linear model. Further, transformations of continuous variables (such as polynomials) can be added for additional flexibility. Alternatively, the researcher might estimate the heterogeneous treatment effects one variable at a time, which may not capture the full heterogeneity pattern but may provide a simple and intuitive summary of different dimensions of heterogeneity. Section 4 illustrates different approaches and interpretations empirically. (See Calonico et al. (2025) for further examples and discussion.)
@@
As with (2.2), and in contrast to (2.3), the coefficients in (2.5) can be obtained by first fitting separate regressions on either side of the cutoff, and then taking appropriate differences. It is also possible to run separate models for different subgroups of discrete variables via (2.2) to recover the estimates in (2.5), but the joint estimation procedure is simple and intuitive and thus appealing for practice. Joint estimation is natural if additional controls are added: following Calonico et al. (2019), it is possible to include some covariates for efficiency gains, while other covariates for heterogeneity analysis, thereby obtaining more efficient estimation and inference for $\kappa(\mathbf{w})$. This mixed covariate adjustment approach augments (2.5) with additional covariates $\mathbf{Z}_{i}$ that are not interacted with $T_{i}$ nor $X_{i}$, but are interacted with $\mathbf{W}_{i}$. That is, one would fit